# MASQUE: A Text-Guided Diffusion-Based Framework for Localized and Customized Adversarial Makeup

### v1

@@ 2.2 Anti-Facial Recognition
where $\mathrm{AFR}:\mathcal{X}\rightarrow\mathcal{X}$ denotes the AFR perturbation function, $\Delta$ is a similarity metric that measures the visual distortion of the AFR-perturbed image $\mathrm{AFR}(x)$ with reference to the original image $x$, and $\gamma>0$ is a threshold parameter reflecting the distortion upper bound the user can tolerate.
Dodging Success. We consider benign users who seek privacy protection through adversarial-based AFR before sharing their facial images online. A desirable AFR technique should achieve a high dodging success rate (DSR) against unauthorized facial recognition, as defined by the optimization objective in Equation LABEL:eq:adversarial_AFR. For clarity, we refer to these users as defenders, as they apply AFR techniques to modify their facial images with adversarial perturbations. Since the defender typically lacks precise knowledge of the FR model employed by the attacker, we consider the black-box scenario when evaluating the DSR of each AFR method.
Visual Quality. From the user’s perspective, images perturbed by AFR should appear natural and closely resemble the original, as captured by the optimization constraint in Equation LABEL:eq:adversarial_AFR. If $\mathrm{AFR}(x)$ has an unsatisfactory image quality, users may be reluctant to share the protected face image online, even if a high DSR is achieved. Given this expectation of high visual quality, earlier privacy-preserving techniques such as face obfuscation or anonymization [20, 29] are not suitable. In comparison, makeup-based AFR methods [26, 30] stand out, as they introduce natural adversarial perturbations with minimal visual distortion. In this work, we further improve the performance of generative makeup-based AFR, enhancing both DSR and visual quality.
Figure 1: Each row presents the reference image with a text makeup prompt, the original image, followed by the generations of each AFR method. Methods marked with $\dagger$ require only a text prompt for makeup application and do not rely on a reference image. The yellow text indicates the cosine similarity score. Below the images, we compare ISR and DSR averaged across four FR models and three makeup styles.
2.3 Why Dodging instead of Impersonation?
In contrast to the untargeted dodging objective, several state-of-the-art studies [37, 12, 26, 30] have focused on improving AFR performance in targeted impersonation scenarios. Impersonation-oriented methods aim to adversarially modify a user’s facial images to mislead a FR model into recognizing them as a specific target identity. Let $\mathrm{AFR}_{\mathrm{I}}:\mathcal{X}\rightarrow\mathcal{X}$ be the perturbation function for an impersonation-oriented method. In existing literature, the impersonation success rate (ISR) is typically measured by the following metric:
$\text{ISR}:=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\mathds{1}\big{\{}\mathrm{FR}\big{(}\mathrm{AFR}_{\mathrm{I}}(x)\big{)}=y_{t}\},$ (3)
where $y_{t}$ denotes the target identity satisfying $y_{t}\neq y$. Below, we further detail why impersonation is not suitable for AFR.
Misaligned Goal. From the defender’s perspective, impersonation and dodging serve different goals, as reflected in their distinct evaluation metrics, ISR and DSR. As shown by [42], high ISR does not guarantee high DSR. If the target identity $y_{t}$ is not included in the attacker’s model (e.g., using facial verification with only a single image of the victim identity $y$), optimizing for higher ISR offers little benefit due to misaligned objectives. As we will illustrate in Section 3.1, prior impersonation-oriented AFR methods can only achieve weak privacy protection when evaluated under dodging scenarios. Therefore, we argue that effective privacy protection should prioritize untargeted dodging success.
Targeted Misuse. Impersonation increases the risk of targeted misuse, raising serious ethical and fairness concerns. Unlike dodging, impersonation-based approaches allow users to control the target identity they are mistaken for, making them more susceptible to abuse. This ability can be exploited for malicious purposes, such as falsely associating one’s actions with another individual. For instance, if a user $y$ intentionally applies an AFR tool to impersonate a target $y_{t}$ and engages in illegal activity, FR systems may misidentify $y_{t}$ as the suspect. This deliberate identity manipulation poses significant ethical and legal risks [34].
By contrast, dodging-oriented AFR serves as a better solution for facial privacy protection. Instead of redirecting recognition to a specifically chosen identity, dodging causes misclassification in an uncontrolled manner, typically as an unknown or incorrect identity. This lack of control makes dodging significantly less susceptible to intentional misuse, as users cannot deliberately impersonate or frame a different party. While dodging still prevents correct identification, it does so without assigning a specific false identity, reducing the risk of targeted deception or malicious intent.
Dodging External ID Guidance Local Prompt-Following
TIP-IM [37] n/a n/a
AMT-GAN [12] image low
C2P [26] text medium
DiffAM [30] image medium
MASQUE (ours) text high
Table 1: Comparisons of SOTA AFR methods for key features in terms of both privacy protection and visual quality.
3 Challenges of Existing AFR Methods
In this section, we present preliminary results to highlight key limitations of previous work, further motivating the design of our approach, which will be detailed in Section 4.
Preliminary Results. Figure 1 shows qualitative comparisons of three makeup styles, evaluating several prior AFR methods, including a noise-based method, TIP-IM [37], and three state-of-the-art makeup-based approaches [12, 26, 30]. For each AFR method, we report the metrics across different black-box facial verification models averaged across $100$ randomly selected images from CelebA-HQ [14]. To ensure a fair comparison with image-guided AFR methods, we either select the corresponding makeup image from a benchmark facial makeup dataset [16] or generate one by inpainting the makeup on a non-makeup image (see Section 5.1 for more details). Table 1 summarizes the key feature comparisons of state-of-the-art AFR methods, including ours.
3.1 Weak Protection under Dodging
As discussed in Section 2.3, an ideal AFR should account for dodging scenarios, to ensure privacy without dependence on a target identity. However, despite its ethical advantages, dodging remains an underexplored and underdeveloped area in makeup-based AFR. Existing methods that perform well under impersonation settings often exhibit a significant drop in DSR, as shown in Figure 1. This highlights a fundamental gap: current techniques are not optimized for dodging, leaving privacy protection less reliable when a user seeks to avoid recognition rather than impersonate another identity. In addition, while some methods like [26, 7] introduce a form of dodging, they still rely on a target identity. This reliance contradicts the very goal of dodging-based AFR, which should function without any identity substitution. Thus, a key challenge in AFR research is developing a robust dodging-based approach that provides strong protection while avoiding the ethical risks associated with impersonation.
3.2 Non-Localized Face Modification
Noise-based approaches like TIP-IM generate global perturbations, which fail to localize the editings. Makeup-based AFR embeds perturbations within makeup to achieve a more natural disguise, but many GAN-based approaches, including AdvMakeup, AMT-GAN, and C2P, introduce global artifacts that alter non-facial regions. This lack of precise localization is particularly problematic when users seek to protect only their facial identity while preserving the background. As shown in Figure 1, this issue is especially pronounced in complex backgrounds, where perturbations should remain confined to facial regions but instead extend beyond. In contrast, the diffusion-based method DiffAM enables better localization, effectively applying makeup perturbations from a reference image. However, its reliance on reference images limits flexibility, particularly when suitable references are unavailable. This highlights the need for methods that achieve both precise localization and greater flexibility.
3.3 Limited Prompt-Following Ability
Makeup-based AFR methods aim to modify facial appearance to prevent accurate identification by FR systems. However, existing methods offer limited user control, making customization and usability a key challenge. Broadly, makeup-based AFR methods fall into two categories:
  • Image-reference-based AFR transfers makeup styles from a reference image. While they aim for accurate replication, they struggle with reliability and flexibility. As shown in Figure 1, early GAN-based methods like AMT-GAN fail to transfer makeup consistently, while DiffAM, despite improvements, remains restricted. It only applies makeup to three fixed regions—skin, mouth, and eyes—preventing transfer to other areas like colored eyebrows. Additionally, these methods require separate fine-tuning for each reference, limiting efficiency.
4 Our Methodology: MASQUE
To address the key challenges discussed in Section 3, we propose MASQUE, a method designed to disrupt FR models with localized adversarial makeup while ensuring no external identity is introduced. Figure 2 illustrates the pipeline of MASQUE with its pseudocode in Appendix D. The makeup is guided by a user-defined text prompt $p^{*}$, refined via cross-attention fusion with a mask. Pairwise adversarial guidance is further introduced to ensure these perturbations mislead FR systems without compromising visual fidelity.
Our method leverages the state-of-the-art Stable Diffusion [24] adapted to generate high-quality adversarial makeups by iterative denoising a random noise latent $z_{T}$ conditioned on a text embedding $C$. To be more specific, Stable Diffusion is trained to predict the added noise $\epsilon$ via:
$\min_{\theta}\ \mathbb{E}_{{z}_{0},\epsilon\sim\mathcal{N}(0,I),t\sim\mathrm{Unif}(1,T)}\bigg{[}\big{\|}\epsilon-\epsilon_{\theta}(z_{t},t,C)\big{\|}_{2}^{2}\bigg{]}.$
Applying Makeup with Cross Attention. Before applying edits, we first obtain a faithful latent representation of the original image $x$ using null-text inversion [19], which mitigates reconstruction errors common in direct DDIM inversion [28]. Conditioning on an empty prompt to align the forward and reverse diffusion trajectories, ensuring near-perfect reconstruction of the image’s structure and identity.
With this accurate latent representation, we introduce makeup attributes via the text prompt $p^{*}$ by manipulating the cross-attention (CA) layers of the diffusion model [24], which control how spatial features correspond to semantic tokens. At each diffusion step $\tau$, we extract attention maps $A_{\tau}$ (reconstruction) and $A_{\tau}^{*}$ (editing) and blend them: preserving CA values from $A_{\tau}$ for shared tokens to maintain structure, while incorporating values from $A_{\tau}^{*}$ for makeup-specific tokens in $p^{*}$:
$\big{(}\mathrm{Update}(A_{\tau},A_{\tau}^{*})\big{)}_{i,j}:=\begin{cases}(A_{\tau})_{i,j},&\text{if $j$ is in both $p$ and $p^{*}$},\\[6.0pt] (A_{\tau}^{*})_{i,j},&\text{if $j$ is unique to $p^{*}$},\end{cases}$
@@ 4 Our Methodology: MASQUE
where $p$ denotes the original text prompt. The result is $\hat{A}_{\tau}$, a set of mixed CA maps that preserve the original facial layout while steadily introducing adversarial makeup features [9].
@@
Enhancing Semantic Edits and Locality. To ensure precise localization, we generate a mask $\mathcal{M}$ that defines the region for modification. To achieve this, we embed the prompt $p^{*}$ using a Sentence Transformer [23] model and compare it to embeddings of predefined facial regions. The closest match determines the relevant area for the edit. For instance, if the prompt specifies “a face with red lipstick”, the model identifies lips as the target and generates a lip-area mask. This ensures that our adversarial makeup perturbations remain focused on the correct facial region and do not unintentionally alter other parts of the face.
Once the target region $\mathcal{M}$ is determined, we enhance the influence of makeup-related tokens by maximizing their attention within $\mathcal{M}$. This ensures that modifications are reinforced within the intended area, preventing unintended changes elsewhere and maintaining overall image quality [18]. Specifically, we optimize:
$\mathcal{L}_{\mathrm{edit}}=\left(1-\frac{1}{|\mathcal{M}|}\sum_{i\in\mathcal{M}}\frac{(A_{\tau}^{*})_{i,\mathrm{new}}}{(A_{\tau}^{*})_{i,\mathrm{new}}+\sum_{j\in\mathrm{share}}(A_{\tau})_{i,j}}\right)^{2},$
@@
where $(A_{\tau}^{*})_{i,\mathrm{new}}$ is the attention weights assigned to the new makeup tokens at spatial index $i$, while $\sum_{j\in\mathrm{share}}(A_{\tau})_{i,j}$ is the total attention weight of tokens that appear in both the original and makeup prompts. By emphasizing makeup-specific attention in the masked region, this loss term ensures that modifications in $x_{\mathrm{p}}$ occur meaningfully in the intended region while preventing edits from spreading elsewhere.
To further enforce spatial precision, we process the latent through two distinct branches: an edit branch for makeup application and a reconstruction branch to preserve original features. During the backward steps, we explicitly constrain perturbations within the designated area by imposing the following conditions:
$z_{\text{edit}}=\mathcal{M}\cdot z_{\text{edit}}+(1-\mathcal{M})\cdot z_{\text{rec}},$
where $z_{\text{edit}}$ represents the latent from the edit branch, and $z_{\text{rec}}$ denotes the latent from the reconstruction branch.
@@
While makeup edits ensure semantic plausibility, the core adversarial objective is to disrupt FR models. Here lies our novel contribution: a pairwise adversarial guidance that uses a guide image of the same identity to achieve robust identity confusion without relying on an external target.
@@
Identification Clean $10.00$ $13.00$ $5.00$ $40.00$ $13.00$ $18.00$ $18.00$ $25.00$ $17.75$
TIP-IM $62.00$ $86.00$ $65.00$ $74.00$ $57.00$ $73.00$ $52.00$ $62.00$ $66.38$
AMT-GAN $61.22$ $48.33$ $50.67$ $57.33$ $16.00$ $24.00$ $22.33$ $32.33$ $39.03$
DiffAM $54.00$ $57.67$ $59.00$ $74.33$ $49.67$ $51.67$ $54.00$ $70.67$ $58.88$
C2P (I) $30.67$ $38.33$ $22.00$ $56.67$ $18.00$ $20.00$ $21.67$ $32.00$ $29.92$
C2P (D) $74.67$ $76.33$ $56.33$ $77.33$ $18.33$ $19.67$ $20.33$ $30.33$ $46.67$
MASQUE ($G=0$) $92.33$ $95.67$ $61.33$ $87.00$ $69.67$ $76.33$ $45.33$ $81.67$ $76.17$
MASQUE ($G=1$) $\underline{98.00}$ $\underline{98.33}$ $\underline{77.67}$ $\underline{94.00}$ $\underline{84.33}$ $\underline{89.00}$ $\underline{61.00}$ $\underline{91.00}$ $\underline{85.50}$
Verification Clean $5.00$ $5.00$ $4.00$ $10.00$ $13.00$ $17.00$ $15.00$ $30.00$ $12.38$
TIP-IM $44.00$ $60.00$ $51.00$ $40.00$ $46.00$ $51.00$ $59.00$ $46.00$ $49.63$
AMT-GAN $40.00$ $32.00$ $51.67$ $25.67$ $14.00$ $26.33$ $26.67$ $40.67$ $32.13$
DiffAM $31.00$ $26.00$ $54.00$ $43.33$ $46.33$ $50.67$ $56.00$ $82.67$ $48.75$
C2P (I) $12.33$ $11.33$ $14.33$ $21.33$ $16.33$ $18.67$ $19.33$ $37.33$ $18.88$
C2P (D) $52.33$ $46.67$ $40.33$ $41.67$ $17.00$ $19.67$ $18.33$ $38.00$ $34.25$
MASQUE ($G=0$) $89.33$ $90.67$ $57.67$ $76.33$ $62.33$ $71.33$ $42.33$ $86.00$ $72.00$
MASQUE ($G=1$) $\underline{96.00}$ $\underline{96.67}$ $\underline{75.67}$ $\underline{79.33}$ $\underline{82.33}$ $\underline{91.67}$ $\underline{63.33}$ $\underline{94.33}$ $\underline{84.92}$
Table 2: Dodging success rate (%) under black-box facial identification and verification settings. For each column, the other three FR systems are used as surrogates to generate the protected faces, and the best result is underlined. Here, $G$ denotes the number of guide images.
Pariwise Adversarial Guidance. Previous makeup-based AFR methods often target another identity, compromising privacy and limiting applicability in dodging scenarios. By contrast, our approach leverages a pair $(x,\tilde{x})$ of face images from the same individual, where $\tilde{x}$ serves as the guide image. This strategy highlights a significant issue with naively using the distance from the original image as the adversarial loss. Note that in the standard diffusion denoising process, the objective is to generate images similar to the original. If we merely maximize the distance from the original image as an adversarial loss, it can potentially create conflicting objectives, leading to an unstable performance in both image quality and adversarial effectiveness (see Tables 2-3 in Appendix B for supporting evidence and comparison results with self-augmented guide images).
Adversarial and Image Quality Trade-off. We introduce adversarial perturbations during the later stage of the diffusion process, ensuring coarse structure remains intact while subtly altering identity-specific features. To balance adversarial potency with visual fidelity, we incorporate perceptual similarity constraints alongside a cosine similarity (CoSi) measure that captures adversarial effectiveness:
$\mathcal{L}_{\mathrm{adv}}=\lambda_{\mathrm{CoSi}}\cdot\mathrm{CoSi}(x_{\mathrm{p}},\tilde{x})+\lambda_{\mathrm{LPIPS}}\cdot\mathrm{LPIPS}(x_{\mathrm{p}},x),$
where $x_{\mathrm{p}}$ denotes the protected image. $\mathrm{CoSi}$ ensures that the adversarial perturbations sufficiently diverge from recognizable identity features, while $\mathrm{LPIPS}$ [39] maintains perceptual and structural fidelity, respectively. The parameters $\lambda_{\mathrm{CoSi}},\lambda_{\mathrm{LPIPS}}$ allow fine-tuning of this trade-off. We conduct ablation studies to understand the sensitivity of these hyperparameters and how to set them in MASQUE for better trade-offs (see Section 5.3 and Appendix B).
5 Experiments
In this section, we conduct comprehensive experiments to study the effectiveness of MASQUE with comparisons to prior AFR methods in both privacy protection and visual aspects. Appendices B and C present additional experimental results and visualizations for more general makeup prompts.
5.1 Experimental Settings

### v2

@@ 4.2 Detailed Design
where $p$ denotes the original text prompt. The result is $\hat{A}_{\tau}$, a set of mixed CA maps that preserve the original facial layout while steadily introducing adversarial makeup features (Hertz et al., 2022).
@@
Enhancing Semantic Edits and Locality. To ensure precise localization, we generate a mask $\mathcal{M}$ that defines the region for modification. To achieve this, we embed the prompt $p^{*}$ using a Sentence Transformer (Reimers & Gurevych, 2019) model and compare it to embeddings of predefined facial regions. The closest match determines the relevant area for the edit. For instance, if the prompt specifies “a face with red lipstick”, the model identifies lips as the target and generates a lip-area mask. This ensures that our adversarial makeup perturbations remain focused on the correct facial region and do not unintentionally alter other parts of the face. Once the target region $\mathcal{M}$ is determined, we enhance the influence of makeup-related tokens by maximizing their attention within $\mathcal{M}$. This ensures that modifications are reinforced within the intended area, preventing unintended changes elsewhere and maintaining overall image quality (Mao et al., 2023). Specifically, we optimize:
$\mathcal{L}_{\mathrm{edit}}=\left(1-\frac{1}{|\mathcal{M}|}\sum_{i\in\mathcal{M}}\frac{(A_{\tau}^{*})_{i,\mathrm{new}}}{(A_{\tau}^{*})_{i,\mathrm{new}}+\sum_{j\in\mathrm{share}}(A_{\tau})_{i,j}}\right)^{2},$ (5)
@@
where $(A_{\tau}^{*})_{i,\mathrm{new}}$ is the attention weights assigned to the new makeup tokens at spatial index $i$, while $\sum_{j\in\mathrm{share}}(A_{\tau})_{i,j}$ is the total attention weight of tokens that appear in both the original and makeup prompts. By emphasizing makeup-specific attention in the masked region, this loss term ensures that modifications in $\bm{x}_{\mathrm{p}}$ occur meaningfully in the intended region while preventing edits from spreading elsewhere. To enforce spatial precision, we process latents through two distinct branches: an edit branch for makeup application and a reconstruction branch to preserve original features. During the backward steps, we explicitly constrain perturbations within the designated area by imposing:
$\bm{z}_{\text{edit}}=\mathcal{M}\cdot\bm{z}_{\text{edit}}+(1-\mathcal{M})\cdot\bm{z}_{\text{rec}},$ (6)
@@
where $\bm{z}_{\text{edit}}$ denotes the latent from the edit branch, and $\bm{z}_{\text{rec}}$ is the latent from the reconstruction branch. While makeup edits ensure semantic plausibility, the core adversarial objective is to disrupt FR models. Therefore, we propose a novel pairwise adversarial guidance loss that uses a guide image of the same identity to achieve robust identity confusion without relying on an external target.
Pariwise Adversarial Guidance. Previous makeup-based AFR methods often target another identity, compromising privacy and limiting applicability in dodging scenarios. By contrast, our approach leverages a pair $(\bm{x},\tilde{\bm{x}})$ of face images from the same individual, where $\tilde{\bm{x}}$ serves as the guide image. This strategy highlights a significant issue with naively using the distance from the original image as the adversarial loss. Note that in the standard diffusion denoising process, the objective is to generate images similar to the original. If we merely maximize the distance from the original image as an adversarial loss, it can potentially create conflicting objectives, leading to an unstable performance in both image quality and adversarial effectiveness. In our evaluations, we test the impact of the number of guide images $G$ on MASQUE’s performance (Appendix C.1) and whether using self-augmentation without a separate guide image ($G=0$) in MASQUE can be an alternative (Appendix C.2).
Adversarial and Image Quality Trade-off. We introduce adversarial perturbations during the later stage of the diffusion process, ensuring the coarse structure remains intact while subtly altering identity-specific features. To balance adversarial potency with visual fidelity, we incorporate perceptual similarity constraints alongside a cosine similarity (CoSi) measure:
$\mathcal{L}_{\mathrm{adv}}=\lambda_{\mathrm{CoSi}}\cdot\mathrm{CoSi}(\bm{x}_{\mathrm{p}},\tilde{\bm{x}})+\lambda_{\mathrm{LPIPS}}\cdot\mathrm{LPIPS}(\bm{x}_{\mathrm{p}},\bm{x}),$ (7)
where $\bm{x}_{\mathrm{p}},\tilde{\bm{x}}$ refer to the protected and guide images, and $\lambda_{\mathrm{CoSi}},\lambda_{\mathrm{LPIPS}}$ are trade-off parameters. Here, $\mathrm{CoSi}(\bm{x}_{\mathrm{p}},\tilde{\bm{x}})$ ensures that the perturbations sufficiently diverge from recognizable identity features, while the LPIPS loss (Zhang et al., 2018) maintains perceptual and structural fidelity.
5 Experiments
In this section, we conduct experiments to benchmark the performance of MASQUE in dodging success and visual aspects, with comparisons to state-of-the-art makeup-based AFR methods. Due to space limits, we present additional experiments, ablation studies, and visualizations in Appendices B-E.
Table 2: DSR (%) of various AFR methods under facial identification and verification settings. Here, $G$ denotes the number of guide images used in MASQUE. Each best result is highlighted in bold.
@@ 5 Experiments
Identification Clean 10.0010.0010.00 13.0013.0013.00 5.005.005.00 40.0040.0040.00 13.0013.0013.00 18.0018.0018.00 18.0018.0018.00 25.0025.0025.00 17.7517.7517.75
TIP-IM 62.0062.0062.00 86.0086.0086.00 65.0065.0065.00 74.0074.0074.00 57.0057.0057.00 73.0073.0073.00 52.0052.0052.00 62.0062.0062.00 66.3866.3866.38
AMT-GAN 61.2261.2261.22 48.3348.3348.33 50.6750.6750.67 57.3357.3357.33 16.0016.0016.00 24.0024.0024.00 22.3322.3322.33 32.3332.3332.33 39.0339.0339.03
DiffAM 54.0054.0054.00 57.6757.6757.67 59.0059.0059.00 74.3374.3374.33 49.6749.6749.67 51.6751.6751.67 54.0054.0054.00 70.6770.6770.67 58.8858.8858.88
C2P (I) 30.6730.6730.67 38.3338.3338.33 22.0022.0022.00 56.6756.6756.67 18.0018.0018.00 20.0020.0020.00 21.6721.6721.67 32.0032.0032.00 29.9229.9229.92
C2P (D) 74.6774.6774.67 76.3376.3376.33 56.3356.3356.33 77.3377.3377.33 18.3318.3318.33 19.6719.6719.67 20.3320.3320.33 30.3330.3330.33 46.6746.6746.67
MASQUE ($G=0$) 92.3392.3392.33 95.6795.6795.67 61.3361.3361.33 87.0087.0087.00 69.6769.6769.67 76.3376.3376.33 45.3345.3345.33 81.6781.6781.67 76.1776.1776.17
MASQUE ($G=1$) $\bm{98.00}$ $\bm{98.33}$ $\bm{77.67}$ $\bm{94.00}$ $\bm{84.33}$ $\bm{89.00}$ $\bm{61.00}$ $\bm{91.00}$ $\bm{85.50}$
Verification Clean 5.005.005.00 5.005.005.00 4.004.004.00 10.0010.0010.00 13.0013.0013.00 17.0017.0017.00 15.0015.0015.00 30.0030.0030.00 12.3812.3812.38
TIP-IM 44.0044.0044.00 60.0060.0060.00 51.0051.0051.00 40.0040.0040.00 46.0046.0046.00 51.0051.0051.00 59.0059.0059.00 46.0046.0046.00 49.6349.6349.63
AMT-GAN 40.0040.0040.00 32.0032.0032.00 51.6751.6751.67 25.6725.6725.67 14.0014.0014.00 26.3326.3326.33 26.6726.6726.67 40.6740.6740.67 32.1332.1332.13
DiffAM 31.0031.0031.00 26.0026.0026.00 54.0054.0054.00 43.3343.3343.33 46.3346.3346.33 50.6750.6750.67 56.0056.0056.00 82.6782.6782.67 48.7548.7548.75
C2P (I) 12.3312.3312.33 11.3311.3311.33 14.3314.3314.33 21.3321.3321.33 16.3316.3316.33 18.6718.6718.67 19.3319.3319.33 37.3337.3337.33 18.8818.8818.88
C2P (D) 52.3352.3352.33 46.6746.6746.67 40.3340.3340.33 41.6741.6741.67 17.0017.0017.00 19.6719.6719.67 18.3318.3318.33 38.0038.0038.00 34.2534.2534.25
MASQUE ($G=0$) 89.3389.3389.33 90.6790.6790.67 57.6757.6757.67 76.3376.3376.33 62.3362.3362.33 71.3371.3371.33 42.3342.3342.33 86.0086.0086.00 72.0072.0072.00
MASQUE ($G=1$) $\bm{96.00}$ $\bm{96.67}$ $\bm{75.67}$ $\bm{79.33}$ $\bm{82.33}$ $\bm{91.67}$ $\bm{63.33}$ $\bm{94.33}$ $\bm{84.92}$
5.1 Experimental Setup